# Optimal Quoting under Adverse Selection and Price Reading

### v2

@@ 2.1 Modeling Framework and Notation
We consider a market maker responsible for quoting bid and ask prices on a single asset. The market maker continuously sets quotes and can stream different pricing ladders to different categories of clients. Specifically, clients are divided into $N$ tiers, indexed by the set $\mathcal{N}:=\{1,\ldots,N\}$.
@@
Here, $\Lambda^{n,k,b}$ and $\Lambda^{n,k,a}$ are smooth and decreasing functions, satisfying standard assumptions in market making models à la Avellaneda-Stoikov or Cartea-Jaimungal (see [13]); in particular,
$c^{n,k,b/a}(\delta):=2-\frac{\Lambda^{n,k,b}(\delta)\Lambda^{n,k,b^{\prime\prime}}(\delta)}{\Lambda^{n,k,b^{\prime}}(\delta)^{2}}$
@@ 2.2 The Optimization Problem
In this paper, we instead adopt a stationary infinite-horizon perspective, [^6] In market making models, the focus naturally lies on the optimal behavior far from the terminal time (which is often arbitrary). In the analysis that follows, we use Taylor expansions to capture the first-order impact of informational risk, working directly in an infinite-horizon framework. Performing these expansions in a finite-horizon setting and then taking the limit as $T\to+\infty$ would lead to results with no financial relevance (this is the classical iterated limit problem). and consider the discounted objective:
@@ 3.1 The Hamilton-Jacobi-Bellman Equation
The HJB equation corresponding to the infinite horizon problem (5) is a nonlinear system for the value function $\vartheta:\Delta^{1}\mathbb{Z}\to\mathbb{R}$, given by:
@@ 3.2 The Baseline without Informational Risk
$\tilde{\delta}^{n,k,b/a*}(p):=\text{argmax}_{\delta\in\mathbb{R}}\Lambda^{n,k,b/a}(\delta)\,(\delta-p)=\left(\Lambda^{n,k,b/a}\right)^{-1}\left(-{H^{n,k,b/a}}^{\prime}(p)\right).$
@@ 3.3.3 Optimal Quotes
$\tilde{\mathfrak{d}}^{n,k,b/a*}(q,p^{\cdot,b},p^{\cdot,a})=\tilde{\delta}^{n,k,b/a*}(p^{\cdot,b/a})+\varepsilon g^{n,k,b/a}(q,p^{\cdot,b},p^{\cdot,a})+o(\varepsilon),$
@@
$\displaystyle=$ $\displaystyle\tilde{\delta}^{n,k,b*}(D^{k}_{+}\theta(q)+\varepsilon D^{k}_{+}f(q))+\varepsilon g^{n,k,b}(q,D^{\cdot}_{+}\theta(q)+\varepsilon D^{k}_{+}f(q),D^{\cdot}_{-}\theta(q)+\varepsilon D^{k}_{-}f(q))+o(\varepsilon)$
@@ 4.3 Optimal Quotes
then, as in Section 3.3.3, the first-order effects of informational risk on optimal quotes in feedback form are given by

### v3

@@ 2.1 Modeling Framework and Notation
We consider a market maker responsible for quoting bid and ask prices on a single asset. The market maker continuously sets quotes and can stream different price ladders to different categories of clients. Specifically, clients are divided into $N$ tiers, indexed by the set $\mathcal{N}:=\{1,\ldots,N\}$.
@@
Here, $\Lambda^{n,k,b}$ and $\Lambda^{n,k,a}$ are smooth, positive and decreasing functions, satisfying standard assumptions in market making models à la Avellaneda-Stoikov or Cartea-Jaimungal (see [13]); in particular,
$c^{n,k,b/a}(\delta):=2-\frac{\Lambda^{n,k,b/a}(\delta)\Lambda^{n,k,b/a^{\prime\prime}}(\delta)}{\Lambda^{n,k,b/a^{\prime}}(\delta)^{2}}$
@@ 2.2 The Optimization Problem
In this paper, we instead adopt a stationary infinite-horizon perspective, [^7] In market making models, the focus naturally lies on the optimal behavior far from the terminal time (which is often arbitrary). In the analysis that follows, we use Taylor expansions to capture the first-order impact of informational risk, working directly in an infinite-horizon framework. Performing these expansions in a finite-horizon setting with terminal time $T$ and then taking the limit $T\to+\infty$ would lead to results with no financial relevance (this is the classical iterated limit problem). and consider the discounted objective:
@@ 3.1 The Hamilton-Jacobi-Bellman Equation
The HJB equation corresponding to the infinite-horizon problem (5) is a nonlinear system for the value function $\vartheta:\Delta^{1}\mathbb{Z}\to\mathbb{R}$, given by:
@@ 3.2 The Baseline without Informational Risk
$\tilde{\delta}^{n,k,b/a*}(p):=\underset{\delta\in\mathbb{R}}{\text{argmax}}\ \Lambda^{n,k,b/a}(\delta)\,(\delta-p)=\left(\Lambda^{n,k,b/a}\right)^{-1}\left(-{H^{n,k,b/a}}^{\prime}(p)\right).$
@@ 3.3.3 Optimal Quotes
$\tilde{\mathfrak{d}}^{n,k,b/a*}(q,p^{\cdot,b},p^{\cdot,a})=\tilde{\delta}^{n,k,b/a*}(p^{k,b/a})+\varepsilon g^{n,k,b/a}(q,p^{\cdot,b},p^{\cdot,a})+o(\varepsilon),$
@@
$\displaystyle=$ $\displaystyle\tilde{\delta}^{n,k,b*}(D^{k}_{+}\theta(q)+\varepsilon D^{k}_{+}f(q))+\varepsilon g^{n,k,b}(q,D^{\cdot}_{+}\theta(q),D^{\cdot}_{-}\theta(q))+o(\varepsilon)$
@@ 4.3 Optimal Quotes
then, as in Section 3.3.3, the first-order effects of informational risk on the (approximation of the) optimal quotes in feedback form are given by